# Directly Forecasting Belief for Reinforcement Learning with Delays

### v1

@@ Abstract
Reinforcement learning (RL) with delays is challenging as sensory perceptions lag behind the actual events: the RL agent needs to estimate the real state of its environment based on past observations. State-of-the-art (SOTA) methods typically employ recursive, step-by-step forecasting of states. This can cause the accumulation of compounding errors. To tackle this problem, our novel belief estimation method, named Directly Forecasting Belief Transformer (DFBT), directly forecasts states from observations without incrementally estimating intermediate states step-by-step. We theoretically demonstrate that DFBT greatly reduces compounding errors of existing recursively forecasting methods, yielding stronger performance guarantees. In experiments with D4RL offline datasets, DFBT reduces compounding errors with remarkable prediction accuracy. DFBT’s capability to forecast state sequences also facilitates multi-step bootstrapping, thus greatly improving learning efficiency. On the MuJoCo benchmark, our DFBT-based method substantially outperforms SOTA baselines.
@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive success in various scenarios, including board games (Silver et al., 2016; Schrittwieser et al., 2020), video games (Mnih et al., 2013; Berner et al., 2019) and cyber-physical systems (Wang et al., 2023b, c; Zhan et al., 2024). The timing factor is critical to enable RL in real-world applications, particularly the delays that occur in the interaction between the agent and the environment due to physical distance, computational processing, and signal transmission. Otherwise, delays fundamentally affect the system’s safety (Mahmood et al., 2018), performance (Cao et al., 2020) and efficiency (Hwangbo et al., 2017). Unlike rewards-delays-arising credit assignment issues that have been well studied (Arjona-Medina et al., 2019; Wang et al., 2024), observation-delays and action-delays disrupt the Markovian property of the Markov Decision Process (MDP), posing more challenges in RL. Most of the literature (Kim et al., 2023; Wang et al., 2023a; Wu et al., 2024b) mainly focuses on observation-delays, which has been proved as a superset of action-delays (Katsikopoulos & Engelbrecht, 2003; Nath et al., 2021). Therefore, this work mainly focuses on the RL with delays $\Delta$ in observation: at time step $t$, the agent can not observe the real environment state $s_{t}$ but only the history state $s_{t-\Delta}$ $\Delta$ steps ago.
@@
To improve learning efficiency in augmentation-based methods, belief-based methods (Walsh et al., 2009; Chen et al., 2021a) propose to conduct RL in the original state space $\mathcal{S}$, by estimating the instant unobservable state $s_{t}$ from the information state $x_{t}$, of which the mapping is referred to as belief. The belief function is commonly forecasted recursively (Chen et al., 2021a; Karamzade et al., 2024): for $i=1,\cdots,\Delta$, $s_{t-\Delta+i}$ is estimated by applying approximate dynamic function with the previous state $s_{t-\Delta+i-1}$ and the previous action $a_{t-\Delta+i-1}$. Obtaining estimation of the instant state $s_{t}$, the delayed RL problem is effectively reduced to a delay-free RL problem without enlarging the state space, mitigating the curse of dimensionality. However, this recursive process is evidently affected by the error accumulation of the approximate dynamic function across $\Delta$ steps: the compounding errors grow exponentially with the delays $\Delta$. This fundamental limitation of such recursive methodology for belief forecasting leads to significant performance degradation, especially in environments with long-delayed signals.
@@
This work aims to squarely address the compounding errors in existing belief-based techniques with the recursive strategy, using the direct strategy in sequence modelling. Specifically, we present the Directly Forecasting Belief Transformer (DFBT), a novel directly forecasting belief method by reformulating belief forecasting as a sequence modelling problem. DFBT first represents the information state $x_{t}$ and reward signals $r_{t-\Delta:t-1}$ to $\Delta$ tokens in the form $\{s_{t-\Delta},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta-1}$. Unlike recursively forecasting belief methods that invoke the forward prediction process $\Delta$ times, DFBT simultaneously forecasts the states $s_{t-\Delta+1:t}$ without introducing accumulated errors, effectively addressing the issue of compounding errors. We then integrate DFBT with the Soft Actor-Critic (SAC) method (Haarnoja et al., 2018) in online learning, incorporating multi-step bootstrapping on accurate state predictions generated from DFBT, which further improves learning efficiency. Theoretically, we demonstrate how the compounding errors in recursively forecasting belief affect the RL performance and how it is alleviated by our direct strategy. Empirically, using the D4RL benchmark (Fu et al., 2020), we show that DFBT achieves significantly higher prediction accuracy compared to other belief methods. On the MuJoCo benchmark (Todorov et al., 2012), across various delay settings, we demonstrate that our DFBT-SAC consistently outperforms SOTA augmentation-based and belief-based methods in both learning efficiency and overall performance.
@@
In Section 3, we introduce the delayed RL problem and the concept of belief representation. Section 4 presents our proposed DFBT, which directly forecasts states from delayed observations, avoiding the need for recursive single-step predictions. Additionally, we develop DFBT-SAC by leveraging multi-step bootstrapping on the states predicted by DFBT. Through theoretical analysis in Section 5, we show that DFBT effectively mitigates the compounding errors associated with recursively forecasting belief methods, ensuring superior performance. Finally, in Section 6, we empirically demonstrate the superior prediction accuracy of DFBT on the D4RL datasets and the remarkable efficacy of DFBT-SAC compared to state-of-the-art approaches on MuJoCo across various delay settings. Overall, our key contributions are summarised as follows:
@@
  • We present the Directly Forecasting Belief Transformer (DFBT), a novel directly forecasting belief method that effectively addresses compounding errors in recursively generated belief.
@@
  • We propose DFBT-SAC, a novel delayed RL method which further improves the learning efficiency via multi-step bootstrapping on the DFBT.
@@ 2 Related Works
Different to typical delay-free RL settings, delayed RL is more aligned with real-world complex systems such as robotics (Mahmood et al., 2018; Hwangbo et al., 2017), finance trading (Hasbrouck & Saar, 2013) and transportation (Cao et al., 2020). Depending on the methodology of retrieving the Markovian property, existing delayed RL approaches can be mainly categorized as augmentation-based and belief-based approaches.
Augmentation-based Approach.
@@
Augmentation-based approaches retrieve the Markovian property by augmenting the original state with the sequence of actions within the delays window, optimizing the policy in the resulting augmented MDP to mitigate performance degradation caused by errors in belief representation approximation. Specifically, DIDA (Liotet et al., 2022) learns the delayed policy in the augmented MDP by imitating the behaviours from the delay-free expert policy in the original MDP; DC/AC (Bouteiller et al., 2020) suggest using the delay correction operator to improve the learning efficiency in the augmented MDP; ADRL (Wu et al., 2024b) introduces the auxiliary delayed task with the smaller augmented MDP for bootstrapping the larger augmented MDP; VDPO (Wu et al., 2024a) proposes to learn the delayed policy through the lens of variational inference, improving the learning efficiency via extensive optimisation tools. However, augmentation-based approaches still operate on an increasingly large augmented state space as delays grow, inevitably suffering from the curse of dimensionality and resulting in significant learning inefficiencies.
Belief-based Approach.
@@
The belief-based approach retrieves the Markovian property by forecasting the unobservable state through belief representation, then optimising policy within the original state space. Thus, correspondingly, the prediction accuracy of approximated belief representation is highly related to the overall performance. DATS (Chen et al., 2021a) is the first to propose approximating the belief representation using a Gaussian distribution. Inspired by the world models (Schmidhuber, 1990b, a; Ha & Schmidhuber, 2018), D-Dreamer (Karamzade et al., 2024) employs a recurrent neural network to capture temporal dependencies in the belief through forward prediction in the latent space, enabling adaptation to high-dimensional observations. D-SAC (Liotet et al., 2021) utilises the attention mechanism in the causal transformer to integrate the information across the delays. Unfortunately, the approximation errors of recursively forecasting belief accumulated at each step finally lead to compounding errors, which hinder prediction accuracy and performance.
Time series Forecasting in RL
@@
Time series forecasting plays a critical role in predicting future states in RL, particularly within the context of world model learning (Ha & Schmidhuber, 2018), which simulates the dynamics of the environment, predicting future states and planning effectively without directly interacting with the real environment. World model learning often leverages time series forecasting techniques to capture temporal dependencies, enhancing the agent’s understanding of the environment. Applications of world models include robotics, autonomous driving, and gaming scenarios (Hafner et al., 2019). Additionally, the RL problem can be treated as a sequence modelling problem, which can be effectively solved by the transformer (Vaswani et al., 2017). Directly predicting the action from the desired outcomes (Schmidhuber, 2019), the decision transformer (Chen et al., 2021b) predicts future actions based on historical trajectories and desired outcomes. Similarly, the trajectory transformer (Janner et al., 2021) models trajectories as sequences, enabling policy optimisation by leveraging autoregressive modelling. Therefore, in this paper, we investigate how to leverage the advanced time series forecasting approaches in addressing the issue of compounding errors existing in the recursively forecasting belief. We also present a comprehensive theoretical analysis of how compounding errors seriously degenerate performance as delays are increased.
@@ 3.1 From Delay-free RL to Delayed RL
A delayed RL problem can be formalised as an augmented MDP by applying the augmentation technique (Altman & Nain, 1992; Katsikopoulos & Engelbrecht, 2003) to retrieve the Markovian property. For a delayed RL problem with constant delays $\Delta$, the newly formed MDP is represented as $\langle\mathcal{X},\mathcal{A},\mathcal{P}_{\Delta},\mathcal{R}_{\Delta},\rho_{\Delta},\gamma\rangle$, where $\mathcal{X}:=\mathcal{S}\times\mathcal{A}^{\Delta}$ is the augmented state space, $\mathcal{A}$ is the original action space, the delayed dynamic function is $\mathcal{P}_{\Delta}(x_{t+1}|x_{t},a_{t}):=\mathcal{P}(s_{t-\Delta+1}|s_{t-\Delta},a_{t-\Delta})\delta_{a_{t}}(a^{\prime}_{t})\prod_{i=1}^{\Delta-1}\delta_{a_{t-i}}(a^{\prime}_{t-i})$ where $\delta$ is the Dirac distribution, the delayed reward function is defined as $\mathcal{R}_{\Delta}(x_{t},a_{t}):=\mathop{\mathbb{E}}_{s_{t}\sim b(\cdot|x_{t})}\left[\mathcal{R}(s_{t},a_{t})\right]$ where $b:\mathcal{X}\times\mathcal{S}\rightarrow\left[0,1\right]$ is the belief representation mapping from the augmented state space $\mathcal{X}$ to the original state space $\mathcal{S}$, $\rho_{\Delta}=\rho\prod_{i=1}^{\Delta}\delta_{a_{-i}}$ is the initial augmented state distribution.
@@ 3.2 Belief Representation in Delayed RL
Delayed RL can be viewed as a special form of partially observable RL problem where the observation is contaminated by the noise, instead of being delayed. Therefore, similar to partially observable RL, delayed RL also have the belief representation defined as follows:
@@ 4 Our Approach
In this section, we present the Directly Forecasting Belief Transformer (DFBT), a new directly forecasting belief method. By framing belief forecasting as a sequence modelling problem, DFBT directly predicts the unobservable states. Instead of recursively predicting the next state step-by-step, DFBT can effectively capture the dependency across the delays via the attention mechanism of the transformer model. Therefore, our DFBT can effectively reduce the compounding errors of the recursively forecasting belief, especially in the long delays. Specifically, as summarised in Figure 1, we train the DFBT using a pre-collected trajectory dataset, later deploying the trained DFBT in the environment with delays to reconstruct the delay-free environment for training. Furthermore, we integrate multi-step bootstrapping with the predicted states of DFBT to improve learning efficiency.
@@ 4.1 Directly Forecasting Belief
The belief representation learning problem can be viewed as a sequence modelling problem. Given a pre-collected delay-free trajectory $\{s_{i},a_{i},r_{i}\}_{i=0}^{T}$, we sample the sub-trajectory with $\Delta$ timesteps $\{s_{t-\Delta+i},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta}$. We model the past reward signals $r_{t-\Delta:t-1}$ into the augmented state $x_{t}=\{s_{t-\Delta},a_{t-\Delta:t-1}\}$ for considering more dynamic information in belief. Then, we have reformed the representation of the augmented state to $\Delta$ tokens for sequence modeling: $x^{\text{tokens}}_{t}=\{s_{t-\Delta},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta-1}$. The Directly Forecasting Belief Transformer (DFBT) $b_{\theta}$ leverages the transformer architecture (Vaswani et al., 2017), utilising the attention mechanism to effectively capture dependencies across long delays. Inputting with $x^{\text{tokens}}_{t}$, DFBT simultaneously predicts the unobserved $\Delta$ states $\{s_{t-\Delta+i}\}_{i=1}^{\Delta}$ via autoregressive modeling with loss:

### v2

@@ Abstract
Reinforcement learning (RL) with delays is challenging as sensory perceptions lag behind the actual events: the RL agent needs to estimate the real state of its environment based on past observations. State-of-the-art (SOTA) methods typically employ recursive, step-by-step forecasting of states. This can cause the accumulation of compounding errors. To tackle this problem, our novel belief estimation method, named Directly Forecasting Belief Transformer (DFBT), directly forecasts states from observations without incrementally estimating intermediate states step-by-step. We theoretically demonstrate that DFBT greatly reduces compounding errors of existing recursively forecasting methods, yielding stronger performance guarantees. In experiments with D4RL offline datasets, DFBT reduces compounding errors with remarkable prediction accuracy. DFBT’s capability to forecast state sequences also facilitates multi-step bootstrapping, thus greatly improving learning efficiency. On the MuJoCo benchmark, our DFBT-based method substantially outperforms SOTA baselines. Code is available at https://github.com/QingyuanWuNothing/DFBT.
@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive success in various scenarios, including board games (Silver et al., 2016; Schrittwieser et al., 2020), video games (Mnih et al., 2013; Berner et al., 2019) and cyber-physical systems (Wei et al., 2017; Wang et al., 2023b, c; Zhan et al., 2024). The timing factor is critical to enable RL in real-world applications, particularly the delays that occur in the interaction between the agent and the environment due to physical distance, computational processing, and signal transmission. Otherwise, delays fundamentally affect the system’s safety (Mahmood et al., 2018), performance (Cao et al., 2020) and efficiency (Hwangbo et al., 2017). Unlike rewards-delays-arising credit assignment issues that have been well studied (Arjona-Medina et al., 2019; Wang et al., 2024), observation-delays and action-delays disrupt the Markovian property of the Markov Decision Process (MDP), posing more challenges in RL. Most of the literature (Kim et al., 2023; Wang et al., 2023a; Wu et al., 2024b) mainly focuses on observation-delays, which has been proved as a superset of action-delays (Katsikopoulos & Engelbrecht, 2003; Nath et al., 2021). Therefore, this work mainly focuses on the RL with delays $\Delta$ in observation: at time step $t$, the agent can not observe the real environment state $s_{t}$ but only the history state $s_{t-\Delta}$ $\Delta$ steps ago.
@@
To improve learning efficiency in augmentation-based methods, belief-based methods (Walsh et al., 2009; Chen et al., 2021a) propose to conduct RL in the original state space $\mathcal{S}$, by estimating the instant unobservable state $s_{t}$ from the information state $x_{t}$, of which the mapping is referred to as belief. The belief function is commonly forecasted recursively (Chen et al., 2021a; Karamzade et al., 2024): for $i=1,\ldots,\Delta$, $s_{t-\Delta+i}$ is estimated by applying approximate dynamic function with the previous state $s_{t-\Delta+i-1}$ and the previous action $a_{t-\Delta+i-1}$. Obtaining the estimation of the instant state $s_{t}$, the delayed RL problem is effectively reduced to a delay-free RL problem without enlarging the state space, mitigating the curse of dimensionality. However, this recursive process is evidently affected by the error accumulation of the approximate dynamic function across $\Delta$ steps: the compounding errors grow exponentially with the delays $\Delta$. This fundamental limitation of such recursive methodology for belief forecasting leads to significant performance degradation, especially in environments with long-delayed signals.
@@
This work aims to squarely address the compounding errors in existing belief-based techniques with recursive strategy, using a direct strategy in sequence modeling. Specifically, we present the Directly Forecasting Belief Transformer (DFBT), a novel directly forecasting belief method by reformulating belief forecasting as a sequence modeling problem. DFBT first represents the information state $x_{t}$ and reward signals $r_{t-\Delta:t-1}$ to $\Delta$ tokens in the form $\{s_{t-\Delta},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta-1}$. Unlike recursively forecasting belief methods that invoke the forward prediction process $\Delta$ times, DFBT simultaneously forecasts the states $s_{t-\Delta+1:t}$ without introducing accumulated errors, effectively addressing the issue of compounding errors. We then integrate DFBT with the Soft Actor-Critic (SAC) method (Haarnoja et al., 2018) in online learning, incorporating multi-step bootstrapping on accurate state predictions generated from DFBT, which further improves learning efficiency. Theoretically, we demonstrate how the compounding errors in recursively forecasting belief affect the RL performance and how it is alleviated by our direct strategy. Empirically, using the D4RL benchmark (Fu et al., 2020), we show that DFBT achieves significantly higher prediction accuracy compared to other belief methods. On the MuJoCo benchmark (Todorov et al., 2012), across various delays settings, we demonstrate that our DFBT-SAC consistently outperforms SOTA augmentation-based and belief-based methods in both learning efficiency and overall performance.
@@
In Section 3, we introduce the delayed RL problem and the concept of belief representation. Section 4 presents our proposed DFBT, which directly forecasts states from delayed observations, avoiding the need for recursive single-step predictions. Additionally, we develop DFBT-SAC by leveraging multi-step bootstrapping on the states predicted by DFBT. Through theoretical analysis in Section 5, we show that DFBT effectively mitigates the compounding errors associated with recursively forecasting belief methods, ensuring superior performance. Finally, in Section 6, we empirically demonstrate the superior prediction accuracy of DFBT on the D4RL datasets and the remarkable efficacy of DFBT-SAC compared to state-of-the-art approaches on MuJoCo across various delays settings. Overall, our key contributions are summarized as follows:
@@
  • We present Directly Forecasting Belief Transformer (DFBT), a novel directly forecasting belief method that effectively addresses compounding errors in recursively generated belief.
@@
  • We propose DFBT-SAC, a novel delayed RL method that further improves the learning efficiency via multi-step bootstrapping on the DFBT.
@@ 2 Related Works
In classical control theory, delays are modelled with delay–differential equations (DDEs) (Cooke, 1963). A rich toolbox is available for analysing their reachability (Fridman & Shaked, 2003; Xue et al., 2020), stability (Feng et al., 2019), and safety (Xue et al., 2021). These techniques, however, assume fully known dynamics and become impractical for high-dimensional systems. Delayed reinforcement learning (RL) is far closer to real deployments than the traditional delay approach and the delay-free RL setting that dominates the literature. Latencies are intrinsic to robotics (Mahmood et al., 2018; Hwangbo et al., 2017), high-frequency trading (Hasbrouck & Saar, 2013), intelligent transportation (Cao et al., 2020), and many other domains. Depending on the methodology of retrieving the Markovian property, existing delayed RL approaches can be mainly categorized as augmentation-based and belief-based approaches.
Augmentation-based Approaches.
@@
Augmentation-based approaches retrieve the Markovian property by augmenting the original state with the sequence of actions within the delays window, optimizing the policy in the resulting augmented MDP to mitigate performance degradation caused by errors in belief representation approximation. Specifically, DIDA (Liotet et al., 2022) learns the delayed policy in the augmented MDP by imitating the behaviours from the delay-free expert policy in the original MDP; DC/AC (Bouteiller et al., 2020) suggest using the delays correction operator to improve the learning efficiency in the augmented MDP; ADRL (Wu et al., 2024b) introduces the auxiliary delayed task with the smaller augmented MDP for bootstrapping the larger augmented MDP; VDPO (Wu et al., 2024a) proposes to learn the delayed policy through the lens of variational inference, improving the learning efficiency via extensive optimization tools. However, augmentation-based approaches still operate on an increasingly large augmented state space as delays grow, inevitably suffering from the curse of dimensionality and resulting in significant learning inefficiencies.
Belief-based Approaches.
@@
The belief-based approach retrieves the Markovian property by forecasting the unobservable state through belief representation, then optimizing policy within the original state space. Thus, correspondingly, the prediction accuracy of approximated belief representation is highly related to the overall performance. DATS (Chen et al., 2021a) is the first to propose approximating the belief representation using a Gaussian distribution. Inspired by the world models (Schmidhuber, 1990b, a; Ha & Schmidhuber, 2018), D-Dreamer (Karamzade et al., 2024) employ a recurrent neural network to capture temporal dependencies in the belief through forward prediction in the latent space, enabling adaptation to high-dimensional observations. D-SAC (Liotet et al., 2021) utilizes the attention mechanism in the causal transformer to integrate the information across the delays. Unfortunately, the approximation errors of recursively forecasting belief accumulated at each step finally lead to compounding errors, which hinder prediction accuracy and performance.
Time Series Forecasting in RL.
@@
Time series forecasting plays a critical role in predicting future states in RL, particularly within the context of world model learning (Ha & Schmidhuber, 2018), which simulates the dynamics of the environment, predicting future states and planning effectively without directly interacting with the real environment. World model learning often leverages time series forecasting techniques to capture temporal dependencies, enhancing the agent’s understanding of the environment. Applications of world models include robotics, autonomous driving, and gaming scenarios (Hafner et al., 2019). Additionally, the RL problem can be treated as a sequence modeling problem, which can be effectively solved by the transformer (Vaswani et al., 2017). Directly predicting the action from the desired outcomes (Schmidhuber, 2019), decision transformer (Chen et al., 2021b) predict future actions based on historical trajectories and desired outcomes. Similarly, trajectory transformer (Janner et al., 2021) models trajectories as sequences, enabling policy optimization by leveraging autoregressive modeling. Therefore, in this paper, we investigate how to leverage the advanced time series forecasting approaches in addressing the issue of compounding errors existing in the recursively forecasting belief. We also present a comprehensive theoretical analysis of how compounding errors seriously degenerate performance as delays are increased.
@@ 3.1 From Delay-free RL to Delayed RL
A delayed RL problem can be formalized as an augmented MDP by applying the augmentation technique (Altman & Nain, 1992; Katsikopoulos & Engelbrecht, 2003) to retrieve the Markovian property. For a delayed RL problem with constant delays $\Delta$, the newly formed MDP is represented as $\langle\mathcal{X},\mathcal{A},\mathcal{P}_{\Delta},\mathcal{R}_{\Delta},\rho_{\Delta},\gamma\rangle$, where $\mathcal{X}:=\mathcal{S}\times\mathcal{A}^{\Delta}$ is the augmented state space, $\mathcal{A}$ is the original action space, the delayed dynamic function is $\mathcal{P}_{\Delta}(x_{t+1}|x_{t},a_{t}):=\mathcal{P}(s_{t-\Delta+1}|s_{t-\Delta},a_{t-\Delta})\delta_{a_{t}}(a^{\prime}_{t})\prod_{i=1}^{\Delta-1}\delta_{a_{t-i}}(a^{\prime}_{t-i})$ where $\delta$ is the Dirac distribution, the delayed reward function is defined as $\mathcal{R}_{\Delta}(x_{t},a_{t}):=\mathop{\mathbb{E}}_{s_{t}\sim b(\cdot|x_{t})}\left[\mathcal{R}(s_{t},a_{t})\right]$ where $b:\mathcal{X}\times\mathcal{S}\rightarrow\left[0,1\right]$ is the belief representation mapping from the augmented state space $\mathcal{X}$ to the original state space $\mathcal{S}$, $\rho_{\Delta}=\rho\prod_{i=1}^{\Delta}\delta_{a_{-i}}$ is the initial augmented state distribution.
@@ 3.2 Belief Representation in Delayed RL
Delayed RL can be viewed as a special form of a partially observable RL problem where the observation is contaminated by noise, instead of being delayed. Therefore, similar to partially observable RL, delayed RL also have the belief representation defined as follows:
@@ 4 Our Approach
In this section, we present the Directly Forecasting Belief Transformer (DFBT), a new directly forecasting belief method. By framing belief forecasting as a sequence modeling problem, DFBT directly predicts the unobservable states. Instead of recursively predicting the next state step-by-step, DFBT can effectively capture the dependency across the delays via the attention mechanism of the transformer model. Therefore, our DFBT can effectively reduce the compounding errors of the recursively forecasting belief, especially in the long delays. Specifically, as summarized in Figure 1, we train the DFBT using a pre-collected trajectory dataset, latter deploying the trained DFBT in the environment with delays to reconstruct the delay-free environment for training. Furthermore, we integrate multi-step bootstrapping with the predicted states of DFBT to improve learning efficiency.
@@ 4.1 Directly Forecasting Belief
The belief representation learning problem can be viewed as a sequence modeling problem. Given a pre-collected delay-free trajectory $\{s_{i},a_{i},r_{i}\}_{i=0}^{T}$, we sample the sub-trajectory with $\Delta$ timesteps $\{s_{t-\Delta+i},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta}$. We model the past reward signals $r_{t-\Delta:t-1}$ into the augmented state $x_{t}=\{s_{t-\Delta},a_{t-\Delta:t-1}\}$ for considering more dynamic information in belief. Then, we have reformed the representation of the augmented state to $\Delta$ tokens for sequence modeling: $x^{\text{tokens}}_{t}=\{s_{t-\Delta},a_{t-\Delta+i},r_{t-\Delta+i}\}_{i=0}^{\Delta-1}$. The Directly Forecasting Belief Transformer (DFBT) $b_{\theta}$ leverages the transformer architecture (Vaswani et al., 2017), utilizing the attention mechanism to effectively capture dependencies across long delays. Inputting with $x^{\text{tokens}}_{t}$, DFBT simultaneously predicts the unobserved $\Delta$ states $\{s_{t-\Delta+i}\}_{i=1}^{\Delta}$ via autoregressive modeling with loss: